show that the inequalities satisfies for all point x,y∈R

d*(x,y) ≤d(x,y)≤√n d*(x,y)


1
Expert's answer
2021-04-13T23:27:57-0400

Given inequalities is-


"d^*(x,y)\le d(x,y)\le \sqrt{n}d^*(x,y)"


As x and y belongs to the R i.e. "x,y\in R"


"d^*(x,y)\le d(x,y)~~~~-(1)"


According the rolle's theorem -


"d(x,y)\le \sqrt{n}d^*(x,y)~~~~~~-(2)"


From eqn.(1) and eqn.(2) we have-


"d^*(x,y)\le d(x,y)\le \sqrt{n}d^*(x,y)"